Please help with this.
If you mix 8 liters of a 40% solution of acid with some 70% solutions to get a mixture that is a 50% solution. How much of the 70% solution would be used?

Found 2 solutions by Nate, checkley71:
Answer by Nate(3500)

8(4/10) + x(7/10) = (x + 8)(5/10)
32 + 7x = 5x + 40
2x = 8
x = 4

Answer by checkley71(8403)

.4*8+.7X=.5(X+8) OR .32+.7X=.5X+4 OR .2X=3.68 OR X=18.4L OF 70% SOLUTION